I am have a problem simplifying (x^4y^3z)^3(-2x^3y^4) Please help......
Answer by mathchemprofessor(65)

Using the rule,(a^m)^n= a^m*n
(x^4y^3z)^3(-2x^3y^4) =(x^12y^9z^3)(-2x^3y^4)=-2x^(12+3)y^(9+4)z^3
Rule:(a^m)*(a^n)=a^(m+n)
=-2x^15y^13z^3